A ball is launched from ground level with an initial upwards velocity of 20 m/s and an initial horizontal velocity of 30 m/s. How far from its starting position does the ball land assuming the ground is level?
1
Expert's answer
2020-09-21T08:31:11-0400

Assume that vertical velocity is "u,"

horizontal is "v".

The time for the ball to go up and down:


"t=2\cdot\frac{u}{g}"

The distance the ball will travel horizontally is


"R=t\cdot v=\frac{2vu}{g}=122\text{ m}."